OpenStudy (anonymous):

The sum of 2 CONSECUTIVE integers is greater than 73. - Write an inequality to describe the situation. - Solve the inequality. - Find the pair of integers with the least sum.

OpenStudy (anonymous):

help please??

OpenStudy (jdoe0001):

let's pick one number, ok?

OpenStudy (anonymous):

so say one integer is a, then the next consecutive integer is a + 1 their sum is a + a + 1 and this has to be greater than 73 \[a + a + 1 > 73\] \[2a + 1 > 73\] move the 1 to the other side and divide by two \[a > 36\]

OpenStudy (anonymous):

Yes! I got the same answer but what are the 2 numbers?

OpenStudy (anonymous):

37 and the next consecutive 38 :)

OpenStudy (anonymous):

ohhhhhhh , okay. may you help me on another one please?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay thanks. The length of a rectangle is 2 inches more than its width, and its perimeter is no more than 68 inches - write an inequality to describe the situation. - solve the inequality. -what are the greatest possible dimensions of the rectangle.

OpenStudy (anonymous):

okay so firstly declare variables. lets do length = l, width = w perimiter = p (pretty easy) now use the relationships. l is 2 inches more than w, so l = w + 2. perimiter is no more than 68, so \[p \le 68\] now we want to get rid of p from the inequality because we need our answer in terms of the dimensions... l and w. p = 2l + 2w. replace this p with the one in the inequality. \[2l + 2w \le 68\] we already have l in terms of w, so lets replace l \[2(w + 2) + w \le 68\] \[2w + 2 \le 68\] \[w \le 22\]

OpenStudy (anonymous):

no sorry im not sure

OpenStudy (anonymous):

okay..well thanks anyways.